### v1

@@ Abstract
Modern multi-layer networks are commonly stored and analyzed in a local and distributed fashion because of the privacy, ownership, and communication costs. The literature on the model-based statistical methods for community detection based on these data is still limited. This paper proposes a new method for consensus community detection and estimation in a multi-layer stochastic block model using locally stored and computed network data with privacy protection. A novel algorithm named privacy-preserving Distributed Spectral Clustering (ppDSC) is developed. To preserve the edges’ privacy, we adopt the randomized response (RR) mechanism to perturb the network edges, which satisfies the strong notion of differential privacy. The ppDSC algorithm is performed on the squared RR-perturbed adjacency matrices to prevent possible cancellation of communities among different layers. To remove the bias incurred by RR and the squared network matrices, we develop a two-step bias-adjustment procedure. Then we perform eigen-decomposition on the debiased matrices, aggregation of the local eigenvectors using an orthogonal Procrustes transformation, and $k$-means clustering. We provide theoretical analysis on the statistical errors of ppDSC in terms of eigen-vector estimation. In addition, the blessings and curses of network heterogeneity are well-explained by our bounds.
Keywords: Multi-layer stochastic block model, Distributed SVD, Bias-correction, Differential privacy
@@ 1 Introduction
Advances in modern technology have facilitated the collection of multiple network data from various sources (e.g., Facebook, LinkedIn, and Twitter) that present disparate social ties (Kairouz et al., 2021). This data type can be well represented by multi-layer networks (Mucha et al., 2010; Boccaletti et al., 2014; Kivelä et al., 2014; Han et al., 2015), in which the nodes are the entities of interest, and the network layers reflect their relationships across different sources. As a result, the study of multi-layer networks has received remarkable attention recently. Most existing works assume that all network layers are stored in a single central machine, see, e.g., Paul and Chen (2016); Bhattacharyya and Chatterjee (2018); Liu et al. (2018); Paul and Chen (2020); Lei et al. (2020); Arroyo et al. (2021); MacDonald et al. (2022); Lei and Lin (2022); Huang et al. (2022). However, in practice, the multi-layer network is often locally stored because of the private ownership and the communication cost among others (Dwork et al., 2014; Kairouz et al., 2021). For example, consider a multi-layer friendship network obtained from different social platforms, where the nodes are the common users of these platforms, and the layers correspond to different social ties. In this network, neither the users would like others to know what kind of users they are following completely, nor are the platforms willing to share all the network information because of the ownership. Thus, aggregating all network layers into one central machine would be rather difficult in this scenario. To this end, the analysis of the multi-layer networks should be naturally based on a distributed and federated architecture.
@@
In the literature, the multi-layer stochastic block model (multi-layer SBM) (Han et al., 2015; Paul and Chen, 2016; Barbillon et al., 2017) has been popularly used for community detection in multi-layer networks, in which each layer is assumed to share the same unobserved community membership (Greene and Cunningham, 2013; Liu et al., 2013). There are two major methodological challenges to apply the multi-layer SBM for locally stored networks with privacy-preserving constraints. First, because of the privacy budget, the randomized response (RR) mechanism (Warner, 1965; Karwa et al., 2017; Ayed et al., 2020) that simply flips all edges in the network with a certain probability is often adopted to achieve differential privacy (DP) (Dwork et al., 2006, 2014; Mironov, 2017; Dong et al., 2022; Bi and Shen, 2022). Although this method can greatly protect an individual’s privacy by ensuring that no individual significantly influences the algorithm’s output, the RR mechanism would bring bias into the multi-layer SBM. The existing community detection methods (e.g., Paul and Chen (2020); Arroyo et al. (2021); Lei and Lin (2022); Huang et al. (2022)) proposed for the exact multi-layer SBMs cannot be directly applied in this scenario. Therefore, a bias-correction procedure is needed for community detection and estimation in the composite and RR-perturbed multi-layer SBM.
Second, directly sending the RR-perturbed networks, either with or without bias-correction on the local machines to the central server, can bring large communication cost or storage cost on the server, especially for large networks. When the bias-correction is achieved on the local machines, the networks become dense; when it is performed on the central server, it needs to preserve all networks. Moreover, sending non-debiased sparse networks can be also communication expensive when the network is large. Therefore, a communication-efficient algorithm is of importance in our considered setting.
Motivated by the above problems, we propose a privacy-preserving Distributed Spectral Clustering (ppDSC) algorithm for detecting the communities of locally stored multi-layer networks with private preserving. To avoid possible cancellation of the communities among different layers, we adopt the squared adjacency matrix (Lei and Lin, 2022) as the input of ppDSC. Our proposed algorithm consists of four steps given below.
  • Step 1: we conduct RR on the network located in each local machine to achieve DP.
  • Step 2: we perform bias-correction on each local RR-perturbed network and then obtain a bias-corrected network on each machine.
  • Step 3: we perform eigen-decomposition on each bias-corrected network and send the top-eigenvectors of each network to the central server. Due to the orthogonal ambiguity of subspaces, we average the orthogonal transformed local eigenvectors using the orthogonal Procrustes transformation (OPT) (Schönemann, 1966; Cape, 2020; Charisopoulos et al., 2021) on the central server. This method ensures that the transformed local eigen-vectors are aligned with each other.
  • Step 4: we implement $k$-means clustering on the orthogonalized vector on the central server to obtain the estimated communities.
We provide rigorous non-asymptotic analysis of the statistical errors of ppDSC in terms of eigen-vector estimation and clustering. It turns out that both errors consist of three parts: the squared local, heterogeneity-induced and centralized errors. The privacy effect is inherent in the first and third parts.
1.1 Our contributions and related work
The contribution of this paper is multi-fold, summarized as follows.
First, we propose a novel community detection method for multiple networks collected from different sources with privacy concerns. Our algorithm satisfies the strong notion of DP. For network data, two notions of DP are often used (Rastogi et al., 2009); one is edge-DP, considered in our paper; the other one is node-DP, which conceals the inclusion or removal of a particular node with all its adjacent edges. Under the framework of these two notions, DP algorithms have been developed for several problems, including graph statistics releasing (Nissim et al., 2007; Karwa et al., 2011; Imola et al., 2021), synthetic graphs releasing (Qin et al., 2017), and generative graph models estimation (Mir and Wright, 2012; Borgs et al., 2015; Karwa and Slavković, 2016; Chang et al., 2021), among others. However, there are few studies on privacy-preserving community detection using statistical models, e.g., Nguyen et al. (2016); Ji et al. (2019); Hehir et al. (2021); Seif et al. (2022). In particular, to the best of our knowledge, privacy-preserving community detection for multi-layer networks has not been studied yet.
Second, we develop a bias-correction procedure to remove the bias brought by RR in the multi-layer SBM, and theoretically investigate the effect of the bias-correction on community detection. The previous works focused on deriving the information-theoretical limits (Seif et al., 2022) and misclassification rate of community detection in bias-corrected RR-perturbed SBMs (Hehir et al., 2021). However, how the bias-correction procedure can affect the clustering and estimation accuracy is still unclear, which is studied in our paper through rigorous theoretical justification.
@@ 1.1 Our contributions and related work
Third, we develop a communication-efficient algorithm, ppDSC, by which we only need to send the top eigenvectors from the local machines to the central server. The method is inspired by the distributed PCA proposed for other scenarios in the divide-and-conquer paradigm (Fan et al., 2019; Bhaskara and Wijewardena, 2019; Charisopoulos et al., 2021) and iterative paradigm (Garber et al., 2017; Gang et al., 2019; Chen et al., 2021a; Li et al., 2021; Guo et al., 2021). However, we incorporate the perturbation and bias-correction procedure in our algorithm; as a result, it poses great challenges to derive the theories, including the estimation error bounds.
Last but not least, we provide a series of novel bounds and explicitly show how the bounds depend on the privacy parameters. These theoretical results have not yet been established. In the special case without the privacy and local computation demand, the error bounds of our algorithm reduce to that given in Lei and Lin (2022). The bounds also review the effect of RR on community detection. In particular, adding more network edges can lead to larger error bounds than removing edges. Our established bounds can also help to explain the blessings and curses of heterogeneity among the local networks.
@@ 1.2 Organizations
The rest of the paper is organized as follows. Section 2 introduces the preliminaries of the model and DP. Section 3 develops the algorithm for privacy-preserving community detection of locally stored networks. Section 4 studies the theoretical properties of ppDSC. Section 5 and 6 include the simulation and real data experiments. Section 7 concludes the paper.
@@ 2 Problem formulation
In this section, we first present the multi-layer SBMs for modeling the locally stored networks. Next, we introduce the DP and the RR mechanism for privacy protection.
@@ 2.1 The model for locally stored networks
Suppose that $L$ aligned networks with $n$ common nodes are locally stored in $L$ machines, respectively. The networks are binary and symmetric with the adjacency matrix denoted by $A_{l}\in\{0,1\}^{n\times n},l=1,...,L$. We consider the scenario that $A_{l}$’s share common communities but with possibly different edge densities.
To model the above locally stored networks, we assume that $A_{l}$’s are independently generated from the following multi-layer SBMs. The $n$ nodes are assigned to $K$ non-overlapping communities with the community assignment of node $i$ denoted by $g_{i}\in\{1,...,K\}$. Given the community assignments, each entry $A_{l,ij}(i<j)$ of $A_{l}$ is generated independently according to
$A_{l,ij}\sim{\mathrm{Bernoulli}}(B_{l,g_{i}g_{j}}),\quad i<j,\;l=1,...,L,$ (1)
where $B_{l}$’s are link probability matrices, and the diagonal entries $A_{l,ii}$’s are all 0. The heterogeneity among the networks is inherent in $B_{l}$’s, and we will quantify it in the next section. Note that we assume that $A_{l}$’s are generated locally at the machine $L$’s.
Model (1) is the multi-layer SBMs studied in Han et al. (2015); Paul and Chen (2016) and so forth. When the adjacency matrices $A_{l}$’s are generated and stored in the local machines, we consider DP and the RR mechanism for the adjacency matrices for private protection.
@@ 2.2 Differential privacy and randomized response
Though the local networks $A_{l}$’s are not directly shared by all the local machines, potential privacy breaches still exist (Dwork et al., 2017). DP provides a rigorous framework for privacy-preserving data analysis (Dwork et al., 2006). Generally, a DP algorithm pursues that its output appears similar in probability if one individual changes the input data. By definition, DP protects any individual’s privacy from an adversary who has access to the algorithm output and even sometimes the rest of the data. In addition, it can prevent other kind of attacks, say, the re-identification attack, the tracing attack, and the membership attack, among others; see Dwork et al. (2017) for a survey. In the context of networks, we consider the following notion of edge-DP (Rastogi et al., 2009; Karwa et al., 2017).
@@
Edge-DP protects the privacy of network edges by ensuring that the algorithm’s output does not reveal the inclusion or removal of any certain edge in the network. In other words, edge-DP can protect against an adversary who knows the algorithm’s output and the existence or not of edges between all pairs but one pair of nodes.
@@
Edge-DP could be achieved by flipping the network edges with certain probabilities, which is called the randomized response (RR) mechanism (Warner, 1965; Karwa et al., 2017; Ayed et al., 2020). The algorithm outputs a perturbed adjacency matrix $\tilde{A}$ of $A$. Specifically, for each $i<j$, when $A_{ij}=1$, $\tilde{A}_{ij}$ keeps to be 1 with probability $q$ and flips to 0 otherwise. Similarly, when $A_{ij}=0$, ${\tilde{A}}_{ij}$ remains to be 0 with probability $q^{\prime}$ and changes to 1 otherwise. The diagonal entries ${\tilde{A}}_{ii}=0$. Note that $q$ and $q^{\prime}$ can be published publicly. The relationship between $q,q^{\prime}$ and the privacy budget $\epsilon$ is shown in the following proposition.
@@
According to (3), a typical choice of $q$ and $q^{\prime}$ is that
@@
which has been proven to be optimal in maximizing the mutual information between raw and RR-perturbed data (Ayed et al., 2020).
Although RR can achieve edge-DP, one can imagine that the resulting network largely deviates from the original network. In the next section, we will see how we can debias the perturbed networks so as to improve the algorithm’s utility under the multi-layer SBMs.

### v2

@@ Abstract
This paper proposes a new efficient and scalable consensus community detection approach for multi-layer stochastic block models in a federated learning framework, where network layers are locally stored across several machines and have privacy-preserving needs. Specifically, we develop a spectral clustering-based algorithm named privacy-preserving Distributed Spectral Clustering (ppDSC). To reduce the bias incurred by the randomized response (RR) mechanism for achieving differential privacy, we develop a two-step bias adjustment procedure. For federated learning, we perform the eigen-decomposition of each privacy-preserved network locally and then aggregate the local eigenvectors using orthogonal Procrustes transformation on the central server. We establish a novel bound on the misclassification rate of ppDSC. The new bound reveals the asymmetric roles of the two edge-flipping probabilities of the RR in the misclassification rate. Through the bound, we can also find the optimal choices for the flipping probabilities given a fixed privacy budget. Moreover, we show that ppDSC enjoys the same statistical error rate as its centralized counterpart when the number of machines satisfies a polynomial order with the sample size on each local machine and the effective heterogeneity is well controlled.
Keywords: Bias-correction, Differential privacy, Distributed SVD, Multi-layer stochastic block models
@@ 1 Introduction
Advances in modern technology have facilitated the collection of multiple network data from various sources (e.g., Facebook, LinkedIn, and Twitter) that present disparate social ties (Kairouz et al., 2021). This type of data can be well represented by multi-layer networks (Mucha et al., 2010), where the nodes are the entities of interest, and the network layers reflect their relationships across different local sources. In practice, a multi-layer network is often stored in a distributed manner because of ownership and communication costs, among others (Kairouz et al., 2021). Moreover, the network edges from local sources often contain sensitive and private information (Dwork et al., 2014). Therefore, the analysis of the multi-layer networks should be naturally based on a distributed and federated architecture with additional privacy-preserving guarantees.
@@
In the literature, the multi-layer stochastic block model (multi-layer SBM) (Han et al., 2015; Paul and Chen, 2016; Barbillon et al., 2017) has been popularly used for the community detection task in multi-layer networks, in which each layer is assumed to share the same unobserved community membership. Among the various methods for community detection under multi-layer SBMs, spectral clustering-based methods stand out because of their computational tractability. There are two problems in developing the spectral clustering-based method for multi-layer SBMs in a privacy-preserving and distributed learning paradigm. The first problem is how to obtain theoretically sound statistics for each layer of networks in a privacy-preserving way. The randomized response (RR) mechanism (Warner, 1965; Karwa et al., 2017) that simply flips all edges in the network with a certain probability has been developed to achieve the rigorous differential privacy (DP) notion (Dwork et al., 2006). However, it undoubtedly brings bias into the multi-layer SBMs. Thus, a bias-correction procedure is crucially needed for community detection in the composite and RR-perturbed multi-layer SBMs.
The second problem is how to develop a communication-efficient distributed learning algorithm for detecting communities under multi-layer SBMs. Recent years have seen various spectral clustering-based methods developed for community detection in multi-layer networks, say, the methods based on the aggregated adjacency matrix (e.g., Paul and Chen, 2020; Lei and Lin, 2022) and the tensor-based methods (e.g., Jing et al., 2021). However, these methods are developed in the centralized paradigm, where all network layers are stored on a single central machine. In the federated learning paradigm, directly sending the RR-perturbed networks stored on the local machines to the central server can bring enormous communication costs or storage costs to the server, especially for large networks. Therefore, performing the eigen-decomposition step of spectral clustering on each local machine is a natural way to reduce the communication cost. A follow-up problem is how to combine the eigenvectors to cope with subspaces’ orthogonal ambiguity.
In this paper, we propose a privacy-preserving Distributed Spectral Clustering (ppDSC) algorithm with only one-round of communication for detecting the communities of locally stored multi-layer network. In particular, we assume that the layers of an $L$-layer network are distributed across $m\ (m\leq L)$ machines. In our algorithm, each machine first performs a carefully designed two-step bias-correction procedure on the squared RR-perturbed local networks, where we adopt the squared matrix to avoid possible cancellation of communities among different layers (Lei and Lin, 2022). Then, the eigen-decomposition is performed locally. Finally, the central server collects the local eigenspaces and conducts the spectral clustering using the average of orthogonal Procrustes transformed (OPT) (Schönemann, 1966; Cape, 2020; Charisopoulos et al., 2021) eigenspaces.
The main contributions of this paper are summarized below. First, we propose a one-shot-based distributed spectral clustering algorithm for detecting the communities of RR-perturbed multi-layer SBMs. In particular, we develop a debiased estimator to accommodate the noise incurred by the RR mechanism and the square transformation. The debiased estimator can consistently estimate the original population matrix, which leads to smaller eigenvector and misclassification error bounds than the non-debiased estimator (with respect to its population matrix or the original population matrix). To the best of our knowledge, the theoretical examination of the bias-correction is rarely addressed in existing literature. We propose to aggregate the local eigenspaces instead of the debiased matrices for communication efficiency. Due to orthogonal ambiguity, we propose aligning the local eigenspaces with OPT (Cape (2020); Charisopoulos et al. (2021)). It is worth mentioning that the algorithm is general because it includes the special cases where the number of local machines is one or the privacy-preserving constraint disappears.
Second, we provide novel bounds of the distributed algorithm ppDSC regarding the eigenvector estimation and clustering, which explicitly show how the misclassification rate depends on the privacy parameters. In particular, when the privacy budget in DP increases logarithmically with the number of nodes, the error bounds coincide with those without the privacy-preserving constraint. Two interesting findings are observed from the privacy-aware bounds. First, without a pre-specified privacy budget, perturbing the network entries from 0 to 1 is more harmful than perturbing the network from 1 to 0 to the upper error bound. Second, with a given privacy budget, from our theoretical bounds, one can find the optimal edge-flipping probabilities in the RR mechanism. To the best of our knowledge, these new theoretical results for RR-perturbed DP algorithms have not been investigated in the SBM literature. In addition, the conditions for the bounds reveal that denser networks can allow a more strict privacy-preserving level, i.e., a smaller privacy budget.
Last but not least, the bounds show how local computation and heterogeneity affect the performance of ppDSC. Note that in the regression set-up, the one-shot-based distributed algorithm can attain the centralized bound when the number of machines satisfies a polynomial order of the sample size on each local machine; see Zhang et al. (2013); Lee et al. (2017), among many others. Similarly, we show that the error bound for our ppDSC decomposes into a term that corresponds to the centralized error and two remainder terms corresponding to the squared local error and the effective heterogeneity. The two remaining terms are negligible when the number of machines satisfies a condition similar to the one in the regression setting, and the effective heterogeneity is not severe. Then, the estimator converges at the same rate as a centralized estimator. Our work is the first to derive an error bound for a privacy-preserving and distributed learning algorithm in multi-layer SBMs.
1.1 Related work
Recent years have seen increasing work on multi-layer SBMs including Han et al. (2015); Paul and Chen (2016); Lei et al. (2020); Bhaskara and Wijewardena (2019); Paul and Chen (2020); Arroyo et al. (2021); Lei and Lin (2022); Jing et al. (2021), among many others. A common assumption in these works is that all network layers are centrally stored without considering privacy concerns. As a result, the community detection is performed directly on the original data on the central server. Different from the aforementioned works, we consider the federated learning setting in which the network layers are stored locally on each machine with privacy preservation need. To improve communication efficiency, we propose to work with the locally computed eigenvectors instead of the original data, which results in both methodological and theoretical differences between our methods and the aforementioned works.
There have been various DP algorithms developed for analyzing network data, including graph statistics releasing (Nissim et al., 2007; Karwa et al., 2011; Imola et al., 2021), synthetic graphs releasing (Qin et al., 2017), generative graph models estimation (Mir and Wright, 2012; Borgs et al., 2015; Karwa and Slavković, 2016; Chang et al., 2021), community detection (Nguyen et al., 2016; Ji et al., 2019) among others. These algorithms are not developed based on SBMs. Under the single SBM, two recent works Hehir et al. (2021) and Mohamed et al. (2022) studied respectively the misclassification rate and information-theoretical limits of community detection under the RR mechanism. The differences between our work and these two works are as follows. First, we study the multi-layer SBM while they studied the single SBM. Second, we derive several new privacy-aware bounds that reflect the asymmetric roles of the two edge-flipping probabilities in RR and indicate the best choice of the edge-flipping probabilities given a fixed privacy budget. These observations are new to the literature. In addition, we theoretically show the effect of bias-correction, which was previously unclear.
@@ 1.1 Related work
Driven by large decentralized datasets, there has been increasing work on distributed algorithms for PCA, including the divide-and-conquer paradigm (Fan et al., 2019; Bhaskara and Wijewardena, 2019; Charisopoulos et al., 2021) and iterative paradigm (Garber et al., 2017; Gang et al., 2019; Chen et al., 2021; Guo et al., 2024). However, we work with privacy-preserved network data and incorporate the perturbation and bias-correction procedure in our algorithm; as a result, it poses great challenges to derive the statistical theories, including the estimation error bounds.
1.2 Notation
We use $\lambda_{\min}(A)$ to denote the smallest non-zero eigenvalue of a given matrix $A$. Moreover, we use the projection distance $\mathrm{dist}(U,U^{\prime}):=\|UU^{\intercal}-U^{\prime}U^{\prime{\intercal}}\|_{2}$ to measure the subspace distance between any pair of orthogonal basis $U,U^{\prime}\in\mathbb{R}^{n\times K}$. $\|A\|_{2}$ denotes the spectral norm of the matrix $A$ or the Euclidean norm of vector $A$ and $\|A\|_{\tiny F}$ denotes the Frobenius norm of matrix $A$. In addition, we use the following standard notation for asymptotics. We write $f(n)\asymp g(n)$ or $f(n)=\Theta(g(n))$ if $cg(n)\leq f(n)\leq Cg(n)$ for some constants $0<c<C<\infty$, $f(n)\lesssim g(n)$ or $f(n)=O(g(n))$ if $f(n)\leq Cg(n)$ for some constant $C<\infty$, and $f(n)\gtrsim g(n)$ or $f(n)=\Omega(g(n))$ if $f(n)\geq cg(n)$ for some constant $c>0$. Finally, the constants throughout this work may be different from place to place.
@@ 1.3 Organizations
The rest of the paper is organized as follows. Section 2 provides the preliminaries of the model and DP. Section 3 develops the algorithm for privacy-preserving community detection of locally stored networks. Section 4 studies the theoretical properties of ppDSC. Section 5 and 6 include the simulation and real data experiments. Section 7 concludes the paper.
@@ 2 Problem formulation
In this section, we first present the multi-layer SBMs for generating the locally stored networks. Next, we provide the formal definition of the DP and RR mechanisms.
@@ 2.1 The model for locally stored networks
Suppose there are $L$ aligned networks with $n$ common nodes. The networks are binary and symmetric with the adjacency matrix denoted by $A_{l}\in\{0,1\}^{n\times n},\;l\in[L]:=\{1,2,\dots,L\}$. In the multi-layer SBMs (Han et al., 2015; Paul and Chen, 2016), each $A_{l}$ is independently generated as follows. The $n$ nodes are assigned to $K$ non-overlapping communities with the community assignment of node $i$ denoted by $g_{i}\in[K]$. Given the community assignments, each entry $A_{l,ij}\ (i<j)$ of $A_{l}$ is generated independently according to
$A_{l,ij}\sim{\mathrm{Bernoulli}}(B_{l,g_{i}g_{j}}),\quad i<j,\;l\in[L],$ (1)
where $B_{l}$’s are connectivity matrices, and the diagonal entries $A_{l,ii}$’s are all 0. The heterogeneity among the networks is inherent in $B_{l}$’s, and we will quantify it in the next section.
We assume $A_{l\in[L]}$’s are locally generated and stored at $m$ different machines. For notational simplicity, we assume each machine evenly owns $s:=L/m$ networks. We use $S_{i}:=\{(i-1)s+1,...,i\cdot s\}$ to denote the index set of networks stored in the $i$th machine.
Given the generation process of $A_{l}$’s, we consider the RR mechanism for the adjacency matrices for private protection in the context of DP (Dwork et al., 2014).
@@ 2.2 Differential privacy and randomized response
Though the local networks $A_{l\in S_{i}}$’s are not directly shared by all the local machines, potential privacy breaches still exist (Dwork et al., 2017). DP provides a rigorous framework for privacy-preserving data analysis (Dwork et al., 2006). Generally, a DP algorithm pursues that its output appears similar in probability if one individual changes the input data. By definition, DP protects any individual’s privacy from an adversary who has access to the algorithm output and even sometimes the rest of the data. In addition, it can prevent other kinds of attacks, say, the re-identification attack, the tracing attack, and the membership attack, among others; see Dwork et al. (2017) for a survey. In the context of networks, we consider the following notion of edge-DP (Rastogi et al., 2009; Karwa et al., 2017).
@@
Edge-DP protects the privacy of network edges by ensuring that the algorithm’s output does not reveal the inclusion or removal of any certain edge in the network. In other words, edge-DP can protect against an adversary who knows the algorithm’s output and the existence or absence of edges between all pairs but one pair of nodes.
@@
Edge-DP could be achieved by flipping the network edges with certain probabilities, which is called the randomized response (RR) mechanism (Warner, 1965; Karwa et al., 2017; Ayed et al., 2020). The algorithm outputs a perturbed adjacency matrix $\tilde{A}$ of $A$. Specifically, for each $i<j$, when $A_{ij}=1$, $\tilde{A}_{ij}$ remains 1 with probability $q$ and flips to 0 otherwise. Similarly, when $A_{ij}=0$, ${\tilde{A}}_{ij}$ remains 0 with probability $q^{\prime}$ and changes to 1 otherwise. The diagonal entries ${\tilde{A}}_{ii}=0$. Note that $q$ and $q^{\prime}$ can be published publicly. The relationship between $q,q^{\prime}$ and the privacy budget $\epsilon$ is shown in the following proposition.
@@
According to (3), we can easily derive that for a given $\epsilon$,
[…]
We plot the feasible region of $q$ and $q^{\prime}$ for a given $\epsilon$ in Figure 1, where the formulation of the four boundaries is shown in red. For fixed $\epsilon$, larger $q$ and $q^{\prime}$ will lead to less accuracy loss. Hence, the upper boundaries, i.e., $q^{\prime}=1-\frac{q}{\mathrm{e}^{\epsilon}}$ when $q\leq\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}}$ and $q^{\prime}=\mathrm{e}^{\epsilon}(1-q)$ when $q\geq\frac{\mathrm{e}^{\epsilon}}{1+\mathrm{e}^{\epsilon}}$, are more desirable. A typical choice of $q$ and $q^{\prime}$ is that
@@
which has been proven to be optimal in maximizing the mutual information between raw and RR-perturbed data (Ayed et al., 2020). More importantly, it turns out that this choice is also optimal in the context of this work in the sense that it leads to the smallest misclassification upper bound; see discussions after Theorem 3.
Figure 1: The feasible region (in blue) of $q$ and $q^{\prime}$ under a given privacy budget $\epsilon$.
Although RR can achieve edge-DP, one can imagine that the resulting network largely deviates from the original network. In the next section, we will see how we can de-bias the perturbed networks to improve the algorithm’s utility under the multi-layer SBMs.